OpenStudy (anonymous):

hard question for geniuses Find an equation that generates the series: 1,2,3,3,5,4,7,5 assuming that the first number is at x=1 position, the second is x=2, third x=3 and so on

OpenStudy (anonymous):

\[<a _{n}^{}> _{1} = <1,2,3,3,5,4,7,5,...>\]

OpenStudy (anonymous):

\[\frac{3}{4}n+\frac{1}{2}+\left(\frac{1}{4}n-\frac{1}{2}\right)(-1)^{n+1}\]